# HUG: Hierarchical Urban Gaussian Splatting with Block-Based Reconstruction

### v1

@@ Abstract
As urban 3D scenes become increasingly complex and the demand for high-quality rendering grows, efficient scene reconstruction and rendering techniques become crucial. We present HUG, a novel approach to address inefficiencies in handling large-scale urban environments and intricate details based on 3D Gaussian splatting. Our method optimizes data partitioning and the reconstruction pipeline by incorporating a hierarchical neural Gaussian representation. We employ an enhanced block-based reconstruction pipeline focusing on improving reconstruction quality within each block and reducing the need for redundant training regions around block boundaries. By integrating neural Gaussian representation with a hierarchical architecture, we achieve high-quality scene rendering at a low computational cost. This is demonstrated by our state-of-the-art results on public benchmarks, which prove the effectiveness and advantages in large-scale urban scene representation.
Figure 1: Our method achieves SOTA performance on the MatrixCity dataset. In (a), we show the results of the best-performing baseline method. In (b), we present the results of our method, demonstrating clear improvements in visual quality. For reference, (c) displays the Ground Truth. The quantitative metrics (PSNR, SSIM, LPIPS) of this view shown in (a) and (b) highlight the superiority of our approach in terms of reconstruction quality.
@@ 1 Introduction
Building upon this foundation, we integrate a neural Gaussian representation with an efficient block-wise hierarchical LOD architecture. This combination ensures high-fidelity scene rendering with low computational cost, regardless of distance. By streamlining computational overhead and maintaining high rendering quality, we enable efficient representation of large-scale scenes without compromising performance.
Finally, we demonstrate the effectiveness of our method by achieving State-of-the-Art (SOTA) results on public benchmarks for large-scale scenes as shown in Figure 1.
@@
  • We propose an efficient block-based reconstruction pipeline that optimizes reconstruction quality by cutting unnecessary computational overhead.
  • We propose a hierarchical neural Gaussian representation for in-block training and rendering which enables the efficient representation of large-scale scenes.
  • We achieve SOTA results on public benchmark datasets, demonstrating the advantages of our approach.
@@ 2 Related Work
2.1 Radiance Field Rendering
Radiance field rendering, particularly through Neural Radiance Fields (NeRFs)[20, 3, 2, 19, 25, 24, 32, 34], has made significant strides in 3D scene reconstruction and novel view synthesis. NeRFs represent a scene as a continuous volumetric function, where a deep neural network is trained to map 3D coordinates and viewing directions to radiance and density values. This implicit representation, combined with volumetric rendering, yields highly realistic novel views. However, the method’s dependence on dense sampling and prolonged network inference times often leads to high computational costs and limits its rendering quality. To address this, several approaches have sought to accelerate NeRF-based rendering. Techniques such as InstantNGP [22] leverage multi-resolution hash grids and compact neural networks to achieve substantial speedups while maintaining high visual fidelity. Mip-NeRF 360 [3] extends NeRF’s capabilities to unbounded scenes while maintaining excellent anti-aliasing and high-quality rendering by properly handling multi-scale observations. Similarly, Plenoxels [9] uses sparse voxel grids to efficiently represent the scene’s continuous density field, delivering notable performance improvements.
Figure 2: An overview of our proposed method. (a). The sparse point cloud from COLMAP[27] is used to uniformly divide the scene into blocks. (b). For each camera, visibility masks for all blocks are generated by reprojecting 3D points within blocks into the image space and segmenting corresponding visible regions. (c). Cameras are selected for training block j if the visible area in the mask exceeds a threshold. (d). During training each block, anchors within the visibility mask are selected and optimized with our weighted hierarchical supervision, based on their position and level. Then anchors infer the neural Gaussians used for rendering. (e). All the blocks are seamlessly merged and rendered.
@@ 2.2 Large Scale Scene Reconstruction
Our method addresses these inefficiencies by leveraging the correspondence between geometry and imagery, enabling a targeted reconstruction of the scene’s critical regions. Rather than attempting to reconstruct large surrounding areas, our approach minimizes data redundancy by focusing on the specific regions that require detailed reconstruction. This strategy significantly reduces both computational overhead and memory consumption, offering a more efficient solution for large-scale scene reconstruction.
@@ 2.3 Level of Detail Scene Representation
In the context of learning scene representations, incorporating LOD has become a significant area of active research. Early efforts like Mip-NeRF [2] and NGLOD [30] have utilized hierarchical representations and multi-resolution voxel grids to represent scenes at varying levels of detail. Unlike NeRF-based methods, the LOD handling in 3DGS scene representation is more intuitive. Several Gaussian-based methods[15, 17] achieve the LOD structure by merging the properties of Gaussian volumes after training all the scenes. We take a different approach by integrating LOD construction directly into the training process. Inspired by Octree-GS[26], we use a framework that combines a hierarchical octree structure with Scaffold-GS [18], our approach enables the simultaneous construction of Level-of-Detail (LOD) details and refinement of the scene representation during training, rather than deferring the generation of a compressed LOD model to later stages.
@@ 3 Method
In this section, we introduce HUG, a Hierarchical Urban Gaussian Splatting method designed to efficiently reconstruct and render large-scale urban scenes. An overview of the proposed method is illustrated in Figure 2.
@@ 3.1 Preliminary: 3D Gaussian Splatting
3.2 Focused Scene Partition Strategy
To facilitate efficient reconstruction of large-scale scenes while conserving computational resources, we propose a focused scene partition strategy. This approach partitions the scene into manageable blocks and selects relevant views for each block based on visibility. By doing so, our method enhances reconstruction quality and minimizes computational redundancy.
Data Partition.
Our partitioning approach leverages the sparse point cloud generated by COLMAP [27] as the foundation for dividing the scene. The scene’s bounding box is uniformly divided into multiple blocks indexed by $j$, ensuring controlled block sizes to prevent memory overflow. Each block will be reconstructed independently, and can be processed in parallel across multiple GPUs and machines, allowing for efficient use of computational resources.
@@ 3.2 Focused Scene Partition Strategy
For each image $I_{i}$ and block $j$, we generate a visibility mask $M_{i}^{j}(u,v)$ to indicate which regions of the image correspond to block $j$. The mask is computed by reprojecting the sparse 3D points of block $j$ into the image space using the camera parameters, followed by non-linear classification to segment the visible region:
@@
The mask obtained above is used not only in view selection but also in training stage, detailed in Section 3.3. By concentrating on regions that provide crucial information about the target block. We mitigate the computational waste associated with processing Gaussians that may be discarded in later stages, as observed in other works such as CityGS[17] and VastGaussian [15].
Visibility-Based View Selection.
To select appropriate views for training block $j$, we calculate the total visible area of block $j$ in image $I_{i}$. An image $I_{i}$ is included in the training set for block $j$ if the visible area exceeds a threshold $\tau_{p}$:
$\sum_{(u,v)}M_{i}^{j}(u,v)\geq\tau_{p}.$ (5)
This ensures that only images that significantly cover block $j$ contribute to its reconstruction, thereby reducing computational waste and enhancing efficiency.
3.3 Hierarchical Neural Gaussian Representation
We employ an Octree-based hierarchical reconstruction strategy to optimize each block. Built upon Scaffold-GS[18] and Octree-GS[26], our system has a combination of LOD and Neural Gaussian capabilities. To the best of our knowledge, we are the first to apply a combination of block-trained neural Gaussians and LOD techniques to the reconstruction and rendering of urban scenes. Although GigaGS[5] also use neural anchors during training, these anchors are discarded after training is completed.
Hierarchical Octree Anchor Initialization
Our anchors are initialized at the centers of the nodes of an octree that covers the 3D space. The octree initialization for each block begins with a coarse 3D point cloud $P$, typically obtained from COLMAP[27] or PixSfM[16]. Following the approach in Octree-GS[26], the maximum level of the octree $K$ is determined by $d_{max}$ and $d_{min}$, the 0.95 and 0.05 quantiles of distances between all points in $P$ and the camera centers. It is calculated as follows:
@@ 3.3 Hierarchical Neural Gaussian Representation
Hierarchical Anchor Selection and Optimization
For each training view, a predicted level for each visible anchor is calculated as described in Octree-GS[26]:
@@
where $d$ is the distance between an anchor and this camera center. Only anchors with levels $L\leq\hat{L}$ are then selected for further optimization. As illustrated by the red-marked portion of Figure 3. During optimization, each selected anchor expands into 10 neural Gaussians, each with a explicit offset, neural scaling, rotation, color, and opacity. These neural Gaussians are rendered using same rasterization with 3DGS[12].
In our work, a novel hierarchical training strategy is introduced as shown by the green-marked portion of Figure 3. To enhance hierarchical capabilities and refine details at different levels, each training view will be rendered an additional $K$ times with different LOD levels. We first render anchors with levels $L\leq\hat{L}$ according to Equation 7, the rendered image $R_{full}$ is evaluated with the ground truth image $gt$ using the L1 loss $l_{1}$ and SSIM loss $l_{s}$ losses, weighted by ${\lambda}$ and ${\gamma=1-\lambda}$. Then, we render only the anchors with levels equal to ${K-1}$, and this rendered image $R_{K-1}$ is measured by the two losses mentioned above, but with weight ${\lambda/2^{K}}$ and ${\gamma/2^{1}}$, respectively. For anchors with levels equal to ${K-2}$, the weights will be ${\lambda/2^{K-1}}$ and ${\gamma/2^{2}}$. This process continues until the level reaches 0. Our hierarchical supervision framework incorporates an opacity mask $m^{o}_{L}$, which addresses the presence of extensive transparent regions that arise when rendering higher-level anchors, as illustrated by level 3 in Figure 3 shows. In addition, a weighting factor $\theta$ is introduced to balance the hierarchical component against the full-image content. Accordingly, the hierarchical image-related loss function in our framework can be formulated as follows:
${loss}=$ $\;\lambda\cdot l_{1}(gt,R_{{full}})+\gamma\cdot l_{s}(gt,R_{{full}})$ (8)
@@
This approach is designed to render higher-level anchors with greater detail, while lower-level anchors focus on capturing approximate colors to serve as a foundation for fully rendered images, enhancing hierarchical optimization and improving overall image quality. Our hierarchical anchor supervision strategy can greatly resolve an unavoidable issue in OctreeGS[26] that leads to anomalous rendering. More details are shown in supplementary material.
@@
Anchor Splitting and Pruning
Splitting the Gaussian or anchors based on the gradient is a common operation and was proposed by 3DGS [12]. In our work, anchors with gradients ${\Delta g}$ exceeding a certain threshold $\tau_{g}$ are split into two new anchors with similar properties. For anchors with particularly high gradients, i.e., ${\Delta g}>\beta\tau_{g}$ ($\beta>1$), one generated anchor retains its original level $L$, while the other is assigned a higher level $L+1$, as described in Octree-GS [26]. This operation can enhance the refinement of higher-level details. The splitting is discouraged if specific anchor points are located precisely at the target split node. However, our work found that this restriction degrades performance in detail-rich regions. To improve detail rendering in these areas, we removed this limitation. Additionally, as training progresses, an increasing number of anchors tend to average the loss, eventually causing the gradient threshold to become ineffective. Therefore, we also gradually decrease the gradient threshold for splitting during training, $\tau_{g}(i)=\tau_{g}\cdot\eta^{\lfloor i/M\rfloor}$, where ${i}$ is the current iteration number, $M$ is the decrease interval, and ${\eta}$ is the decrease rate. This greatly improves the splitting efficiency and adaptivity of the training process. To ensure sufficient anchor splitting, we set a percentage threshold ${\tau_{s}}$. If the percentage of anchors meeting the splitting criteria falls below ${\tau_{s}}$, additional anchors with the highest ${\Delta g}$ values within the top ${\tau_{s}}$ will be manually split.
The Octree-GS[26] also graduallys increase the level of each anchor with a gradient exceeding a threshold smaller than $\Delta_{g}$, by adding a property $L_{e}$. We adopt this strategy in our work as well. However, we found that during training, $L_{e}$ can exceed 1 or even reach values greater than 2, indicating that the total LOD level ${K}$, derived from Equation 6, may lack sufficient detail. Therefore, when the average value of the $L_{e}$ is larger than 1, we increase the LOD level of the scene by 1, and reset the $L_{e}$ to lower values. This adjustment allows anchors with higher levels to be generated during the training process.
Pruning anchors or Gaussians based on opacity is a common technique. In addition to this universal pruning strategy, we introduce a visibility-based pruning approach. Since level-related anchor selection is applied in our work, new anchors occasionally appear at incorrect locations with incorrect levels. This results in anchors that retain only their initial properties and are rarely optimized in training views, yet they have to be rendered in test views if they are closer to the test view than to any training views, we show this in the supplementary material. To address this issue, we count the number of times an anchor is selected during training. If this count is below a threshold of ${\epsilon_{c}}$, the anchor is pruned. In addition, we track the accumulated gradient of each anchor and prune those with an accumulated gradient smaller than a threshold of ${\epsilon_{g}}$.
3.4 Local Scene Fusion and Rendering

### v2

@@ Abstract
3DGS is an emerging and increasingly popular technology in the field of novel view synthesis. Its highly realistic rendering quality and real-time rendering capabilities make it promising for various applications. However, when applied to large-scale aerial urban scenes, 3DGS methods suffer from issues such as excessive memory consumption, slow training times, prolonged partitioning processes, and significant degradation in rendering quality due to the increased data volume. To tackle these challenges, we introduce HUG, a novel approach that enhances data partitioning and reconstruction quality by leveraging a hierarchical neural Gaussian representation. We first propose a visibility-based data partitioning method that is simple yet highly efficient, significantly outperforming existing methods in speed. Then, we introduce a novel hierarchical weighted training approach, combined with other optimization strategies, to substantially improve reconstruction quality. Our method achieves state-of-the-art results on one synthetic dataset and four real-world datasets.
Figure 1: Our method achieves the SOTA performance on the MatrixCity dataset. In (a), we show the results of the best-performing baseline method, CityGS [17]. In (b), we present the results of our method, demonstrating clear improvements in visual quality. For reference, (c) displays the Ground Truth. The quantitative metrics for this view, shown in (a) and (b), underscore the superiority of our method in terms of reconstruction quality.
@@ 1 Introduction
Building upon this foundation, we integrate a neural Gaussian representation with an efficient block-wise hierarchical LOD architecture. This combination ensures high-fidelity scene rendering with low computational cost, regardless of distance. By streamlining computational overhead and maintaining high rendering quality, we enable efficient representation of large-scale aerial scenes without compromising performance.
Finally, we demonstrate the effectiveness of our method by achieving State-of-the-Art (SOTA) results on five public benchmarks for large-scale aerial scenes.
@@
  • We propose a simple and effective visibility-based partitioning strategy for large-scale aerial datasets, outperforms existing methods in speed, requiring about 1 minute to process each scene.
  • We introduce a hierarchical neural Gaussian representation tailored for block-based reconstruction of large-scale aerial urban scenes. Our work also presents a novel hierarchical weighted image supervision strategy along with additional optimizations to further enhance reconstruction quality.
  • Our method achieves state-of-the-art results on one synthetic and four real-world datasets, showcasing its advantage over existing approaches.
@@ 2 Related Work
2.1 Neural Rendering
Neural rendering, particularly through Neural Radiance Fields (NeRFs) [20, 3, 2, 19, 25, 24, 32, 34], has made significant strides in 3D scene reconstruction and novel view synthesis. NeRFs represent a scene as a continuous volumetric function, where a deep neural network is trained to map 3D coordinates and viewing directions to radiance and density values. This implicit representation, combined with volumetric rendering, yields highly realistic novel views. However, the method’s dependence on dense sampling and prolonged network inference times often leads to high computational costs and limits its rendering quality. To address this, several approaches have sought to accelerate NeRF-based rendering. Techniques such as InstantNGP [22] leverage multi-resolution hash grids and compact neural networks to achieve substantial speedups while maintaining high visual fidelity. Mip-NeRF 360 [3] extends NeRF’s capabilities to unbounded scenes while maintaining excellent anti-aliasing and high-quality rendering by properly handling multi-scale observations. Similarly, Plenoxels [8] uses sparse voxel grids to efficiently represent the scene’s continuous density field, delivering notable performance improvements.
Figure 2: An overview of our proposed method, HUG. (a). The sparse point cloud from COLMAP [27] is used to uniformly partition the scene into blocks. (b). For each view, visibility masks for all blocks are generated by reprojecting 3D points within blocks into the image space and segmenting corresponding visible regions. (c). Views are assigned to training block $j$ if the visible sparse point cloud exceeds a threshold. (d). During training each block, anchors within the view frustum are selected based on their position and level. These selected anchors infer the neural Gaussians used for rendering and are optimized using our hierarchical weighted image supervision along with other constraints. (e). All trained blocks are seamlessly merged and rendered.
@@ 2.2 Large Scale Scene Reconstruction
Our method overcomes these inefficiencies by utilizing the visibility mask, allowing it to focus on the specific regions that require detailed reconstruction.
@@ 2.3 Level of Detail Scene Representation
In the context of learning scene representations, incorporating LOD has become a significant area of active research. Early efforts like Mip-NeRF [2] and NGLOD [30] have utilized hierarchical representations and multi-resolution voxel grids to represent scenes at varying levels of detail. Unlike NeRF-based methods, the LOD handling in 3DGS scene representation is more intuitive. Several Gaussian-based methods [15, 17] achieve the LOD structure by merging the properties of Gaussian volumes after training all the scenes.
Inspired by Octree-GS [26], we use a framework that combines a hierarchical octree structure with Scaffold-GS [18], our approach enables the simultaneous construction of Level-of-Detail (LOD) details and refinement of the scene representation during training, rather than deferring the generation of a compressed LOD model to later stages.
@@ 3 Method
In this section, we introduce HUG, a Hierarchical Urban Gaussian Splatting method designed for high-quality reconstruction and rendering of large-scale aerial urban scenes. An overview of the proposed method is illustrated in Figure 2.
@@ 3.1 Preliminary: 3D Gaussian Splatting
3.2 Visibility-based Partitioning
To reconstruct large-scale aerial urban scenes, the original 3DGS method must adopt a divide-and-conquer strategy due to memory limitations. We propose a simple yet efficient visibility-based partitioning strategy, which takes only 1 minute on the MatrixCity dataset, whereas CityGS [17] requires approximately 2 hours. Additionally, we integrate an SVM mask derived from visibility, enabling our method to focus more effectively on in-block regions, further enhancing reconstruction quality.
Scene Partitioning.
Our partitioning approach leverages the sparse point cloud $P$ generated by COLMAP [27] as the foundation. The scene is uniformly divided into multiple blocks indexed by $j$, ensuring controlled block sizes to prevent memory overflow. Therefore, the initial sparse point cloud for block $j$ is $P_{j}\subseteq P$.Each block will be reconstructed independently, and can be processed in parallel across multiple GPUs and machines, allowing for efficient use of computational resources.
Visibility-Based View Partitioning.
To determine the appropriate training views for block $j$, we compute the total number of reprojected visible sparse points of block $j$ in image $I_{i}$. An image $I_{i}$ is included in the training set $\mathcal{T}_{j}$ for block $j$ if the number of visible points exceeds a threshold $\tau_{p}$:
$I_{i}\in\mathcal{T}_{j}\quad\text{if}\quad\sum_{p\in P_{j}}\mathbb{I}\left(V(p,I_{i})\right)>\tau_{p}$ (4)
where $\mathbb{I}\left(V(p,I_{i})\right)$ is an indicator function that takes the value 1 if point $p\in P_{j}$ is visible in image $I_{i}$, and 0 otherwise. This selection criterion ensures that only images with sufficient visibility of block $j$ contribute to its reconstruction, reducing redundancy while preserving critical information.
@@ 3.2 Visibility-based Partitioning
For each image $I_{i}$ and block $j$, we generate a visibility mask $M_{i}^{j}(u,v)$ to indicate which regions of the image correspond to block $j$. The mask is computed by reprojecting the sparse 3D points of block $j$ into the image space, followed by non-linear classification to segment the visible region:
@@
The generated mask actively contributes to the training stage by concentrating on regions that provide key information for the target block, thereby reducing computational redundancy and enhancing reconstruction quality within the block.
3.3 Hierarchical Neural Gaussian
We employ an Octree-based hierarchical reconstruction strategy, leveraging neural Gaussians derived from Anchors to optimize each block. Built upon Scaffold-GS [18] and Octree-GS [26], our system has a combination of LOD and Neural Gaussian capabilities. To the best of our knowledge, we are the first to apply a combination of block-trained neural Gaussians and LOD techniques throughout the entire pipeline to the reconstruction and rendering of aerial urban scenes.
Hierarchical Anchor Initialization and Selection
Our anchors are initialized at the centers of the nodes of an octree that covers the 3D space. The octree initialization for each block starts with a coarse 3D point cloud $P_{j}$ corresponding to block $j$, while the complete point cloud $P$ is typically obtained from COLMAP [27] or PixSfM [16]. Inspired by the Octree-GS [26], the maximum level of the octree $K$ is determined by $d_{max}$ and $d_{min}$, the 0.95 and 0.05 quantiles of the distances between all points in $P_{j}$ and all the camera centers. It is calculated as follows:
@@ 3.3 Hierarchical Neural Gaussian
where $d$ is the distance between an anchor and the camera center. Only anchors with levels $L\leq\hat{L}$ are selected for further optimization. As shown by the red-marked portion in Figure 3.
Hierarchical Anchor Optimization
The selected anchors are first decomposed into multiple neural Gaussians, which are then rendered using alpha blending as defined in Equation 3. The resulting image is optimized against the training data using an L1 color loss $l_{1}$ and an SSIM structure loss $l_{s}$, following a similar approach to 3DGS [11] and Octree-GS [26].
In our work, we observe that optimizing hierarchical neural Gaussians using the aforementioned strategy leads to obvious floaters and suboptimal outcomes, as illustrated in Figure 5. To resolve the problems above, a novel hierarchical weighted image supervision is introduced as shown by the green-marked portion of Figure 3.
We begin by rendering anchors with levels $L\leq\hat{L}$ according to Equation 7, the rendered image $R_{full}$ is then compared with the ground truth image $gt$ using the L1 loss $l_{1}$ and SSIM loss $l_{s}$, weighted by ${\lambda}$ and ${\gamma=1-\lambda}$. Next, we render only the anchors with the finest level ${K-1}$, and this rendered image $R_{K-1}$ is evaluated with $l_{1}$ and $l_{s}$, but with weights ${\lambda/2^{K}}$ and ${\gamma/2^{1}}$, respectively. For anchors with levels equal to ${K-2}$, the weights are ${\lambda/2^{K-1}}$ and ${\gamma/2^{2}}$. This process continues until level 0.
Our hierarchical weighted image supervision also incorporates an opacity mask $m^{o}_{L}$, which addresses the presence of extensive transparent regions that arise when rendering higher-level anchors, as shown by level 3 in Figure 3. Additionally, a weighting factor $\theta$ is introduced to balance this component against the traditional training content. Accordingly, the hierarchical loss in our framework can be formulated as follows:
${l_{hier}}=$ $\;\lambda\cdot l_{1}(gt,R_{{full}})+\gamma\cdot l_{s}(gt,R_{{full}})$ (8)
@@
This approach is designed to render higher-level anchors with greater detail, while lower-level anchors focus on capturing approximate colors to serve as a foundation for the fully rendered images. This novel hierarchical weighted image supervision ensures that all anchors remain active during training, enhancing hierarchical optimization and improving overall image quality. Our hierarchical anchor supervision strategy effectively addresses the issue of anomalous rendering in OctreeGS [26], as shown in Figure 5.
@@
Anchor Splitting
Splitting Gaussians or Anchors based on gradients, as proposed in 3DGS [11] and Scaffold-GS [18], involves dividing Gaussians with gradient values $\Delta g$ exceeding a threshold $\tau_{g}$ into two new Gaussians with similar properties.
However, as training progresses, an increasing number of anchors tend to average the loss, causing the gradient threshold $\tau_{g}$ to become less effective. To address this, we propose an innovative dynamic gradient threshold strategy, where the gradient threshold for splitting is gradually decreased during training: $\tau_{g}(i)=\tau_{g}\cdot\eta^{\lfloor i/M\rfloor}$, where ${i}$ is the current iteration number, $M$ is the interval, and ${\eta}$ is the decrease rate.
In the training process, we observed that some anchors have gradients significantly greater than the gradient threshold $\tau_{g}(i)$, indicating that these anchors are unsuitable for remaining in the current LOD. To address this, we propose an innovative dynamic anchor level transition strategy that adaptively refines anchor levels as follow:
$L=\min\left(K-1,L+\left\lfloor\sum(0.01\cdot\mathbb{I}(\Delta g>\beta\cdot\tau_{g}(i)))\right\rfloor\right)$ (9)
where $L$ is the initialized level of an anchor, $\beta$ is a hyperprameter larger than 1, and $\mathbb{I}$ is the indicator function that checks whether the anchor’s gradient at a given training iteration $i$ surpasses $\beta$ times the dynamic gradient threshold. Each time the condition is met, the anchor’s level $L$ is gradually increased by 0.01.
This design allows anchors to dynamically adjust their levels based on local gradient magnitudes, ensuring higher-resolution details in regions where greater precision is needed. By progressively refining anchor levels, our approach enhances the adaptability of the hierarchical anchor representation, leading to improved reconstruction quality.
Anchor Pruning
Pruning anchors or Gaussians based on opacity is a common technique. However, relying solely on opacity may lead to improper optimization of some anchors due to their LOD properties, as illustrated in Figure 5. Previous works [26, 18, 11] often overlooked Gaussians with low visibility counts during pruning.
In our work, we propose an innovative visibility-based anchor pruning strategy specially designed for hierarchical anchors. After every $M$ iterations of training with only distance-based anchor selection, anchors with a visibility count lower than ${\epsilon_{c}}$ will also be pruned.
3.4 Scene Fusion and Rendering